OpenStudy (anonymous):

Calculate the number of phosphorus atoms in 1.7 g of phosphorus. Answer in units of atoms

OpenStudy (lgbasallote):

use dimension analysis \[1.7 \; g\; P \times \frac{1 \; mol \; P}{30.97 \; g\; P} \times \frac{6.022 \times 10^{23} \; atoms \; P}{1 \; mol \; P}\] solve that to get your answer